Fermi National Laboratory
 

Operation Summary for November 13, 2009 - November 16, 2009

 

The Integrated Luminosity for the period from 11/9/09 to 11/16/09 was 51.07 pb-1.

Friday Nov. 13

The day shift began with Operations monitoring the Tevatron (TeV) store 7346 (68.91E30) and the Recycler stash (362.1E10), with the Antiproton Source stacking (24.1E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams.

During the day shift, Operations terminated store 7346 at 9 AM. Cryo system technicians completed TeV cryogenic maintenance at 11:30 AM. TeV personnel completed orbit checks at 1 PM. Operations put the TeV into HEP shot setup at 1:30 PM and established store 7349 at 2:45 PM with an average initial luminosity (AIL) of 271E30. Pbar stacked and the Recycler stashed antiprotons all day except from 2:25 PM to 3:30 PM when maintenance work in Linac and Main Injector inhibited beam. Pbar, NuMI, and MiniBooNE resumed taking beam 3:35 PM.

During the evening shift, the 750 KeV area buncher tripped off; it reset.

Saturday Nov. 14

The midnight shift began with Operations monitoring store 7349 (98.26 E30) and the stash (315.6E10), with Pbar stacking (15.4E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations terminated store 7349 at 2:30 AM. Operations put the TeV into shot setup, but this was interrupted by CDF and D0 collision hall accesses, which lasted until the end of shift.

During the day shift, CDF and D0 were both out of access by 8 AM. Operators conducted TeV orbit checks and then restarted shot setup at 9:30 AM. Operations established store 7352 at 11 AM with an AIL of 264.7E30. There were short downtimes caused by Booster RF stations (BRF13 & 14) and a Linac Klystron RF station (KRF1).

During the evening shift, Operations terminated store 7352 at 10:09 PM and put the TeV into shot setup. Operations established store 7353 at 11:42 PM with an AIL of 260.42E30.

Sunday Nov. 15

The midnight shift began with Operations monitoring store 7353 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

During the midnight shift, at 4 AM a Linac RF station (LRF5) filament failed. Operations called in Linac personnel and they spent the rest of the shift repairing the station.

During the day shift, LRF5 was brought back online and Operations re-established beam by 8:30 AM.

During the evening shift, Operations terminated store 7353 at 7 PM and put the TeV into shot setup. During the shot setup a Main Injector air compressor used to provide air pressure to keep tunnel beam valves open failed. As a result, Recycler beam valves closed and terminated the stash. Operations called in Mechanical techs; they had the Main Injector air compressor fixed by 10:30 PM, however they accessed the Main Injector to troubleshoot problems with beam valve hardware. D0 had accessed their collision hall earlier; they were out of access by 11:10 PM.

Monday Nov. 16

The midnight shift began with mechanical techs accessing the Main Injector, with Operations monitoring the stack, and with NuMI and MiniBooNE waiting for beam.

During the midnight shift, mechanical techs reported that the air pressure for the MI beam valves was not high enough. This was resolved by 12:45 AM. Operations began the MI turn-on. This was halted by a MI sextupole power supply (SEXHPS) failure during turn-on. Power supply personnel had this supply repaired by 3:30 AM. Operations established beam to Pbar, NuMI, and MiniBooNE once the MI access was completed. The Tevatron was in standby until near the end of shift when operators conducted orbit checks and then began a study period.

The Plans for Monday

The plans for today are to stack, stash, and store. MI and mechanical personnel will continue their investigation of why there are different air pressures for opening and closing the MI beam valves.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 11/16/2009    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory